# Evolution of binaries containing a hot subdwarf and a white dwarf to double white dwarfs, and double detonation supernovae with hypervelocity runaway stars

### v1

@@
[…]
11
[…]
Compact binaries containing hot subdwarfs and white dwarfs have the potential to evolve into a variety of explosive transients. These systems could also explain hypervelocity runaway stars such as US 708. We use the detailed binary evolution code MESA to evolve hot subdwarf and white dwarf stars interacting in binaries. We explore their evolution towards double detonation supernovae, helium novae, or double white dwarfs. Our grid of 3120 models maps from initial conditions such as orbital period and masses of hot subdwarf and white dwarf to these outcomes. The minimum amount of helium required to ignite the helium shell that leads to a double detonation supernova in our grid is $\approx 0.05\,\mathrm{M_{\odot}}$, likely too large to produce spectra similar to normal type Ia supernovae, but compatible with inferred helium shell masses from some observed peculiar type I supernovae. We also provide the helium shell masses for our double white dwarf systems, with a maximum He shell mass of $\approx 0.18\,\mathrm{M_{\odot}}$. In our double detonation systems, the orbital velocity of the surviving donor star ranges from $\approx 450\,\mathrm{km\,s^{-1}}$ to $\approx 1000\,\mathrm{km\,s^{-1}}$. Among the surviving donors, we also estimate the runaway velocities of proto-white dwarfs, which have higher runaway velocities than hot subdwarf stars of the same mass. Our grid will provide a first-order estimate of the potential outcomes for the observation of binaries containing hot subdwarfs and white dwarfs from future missions like Gaia, LSST, and LISA.
@@ 1 Introduction
Hot subdwarfs are stripped helium (He) stars that burn He in their core with little to no hydrogen (H) envelope (Heber, 1986, 2016). Studies such as Han et al. (2002, 2003) suggest that binary interactions such as stable or unstable mass transfer plays a major role in the formation of hot subdwarfs. Observations provide evidence that most of the hot subdwarfs are found with binary companions (e.g., Pelisoli et al., 2020). Further, a significant fraction of hot subdwarfs are found in close orbits ($P_{\mathrm{orb}}\sim 10\,\text{days}$) with white dwarfs (WDs), which are likely post common envelope systems (Han et al., 2002, 2003).
@@
Figure 1: A schematic diagram representing possible evolution channels of a hot subdwarf + WD binary to a double white dwarf, a double detonation supernova or a He nova. Double detonation supernovae and He novae result from systems that undergo mass transfer and later experience thermonuclear instability in the He layer on the surface of the WD. The differentiation between double detonation supernova and He nova is based on the He layer’s critical density (see Sec. 2.3). Double white dwarfs result from systems that either evolve in isolation or involve episodes of accretion but do not enter unstable nuclear burning.
@@
Compact binaries containing hot subdwarfs and WDs have gained significant interest following the discovery of a super-Chandrasekhar binary by Maxted et al. (2000). This binary is particularly interesting as a potential Type Ia supernova progenitor (but, see also Ergma et al. 2001 for other outcomes). The number of discoveries of binaries containing hot subdwarfs with WD companions has steadily increased (e.g., Napiwotzki et al., 2004; Geier et al., 2013; Vennes et al., 2012). Large spectroscopic studies have been characterizing binaries containing hot subdwarfs (e.g., Schaffenroth et al., 2022). In addition to those detached binaries, Kupfer et al. (2020a, b) identified binaries in which the hot subdwarf is Roche-lobe filling. Furthermore, some of recent the observed binaries such as CD-30°11223 (Vennes et al., 2012; Geier et al., 2013) and HD 265435 (Pelisoli et al., 2021) are verification binaries for LISA (for more details see Kupfer et al. 2024) and suggested progenitors of double donation supernovae (Kupfer et al., 2022).
@@
Interacting hot subdwarf + WD binaries could potentially result either in an explosion only in the He shell of the WD (a He nova, Kato et al. 2000; Gianninas et al. 2010; Piersanti et al. 2014) or a thermonuclear explosion of a WD (probably as a supernova type Ia, Whelan & Iben 1973; Iben & Tutukov 1984). The thermonuclear explosion of the WD could result from a Chandrasekhar mass explosion (Yoon & Langer, 2003; Wang et al., 2009; Brooks et al., 2016) or from a sub-Chandrasekhar mass double detonation explosion due to He accretion. In the double detonation mechanism, He-accreting WDs undergo the first detonation in the He shell that triggers the second detonation in the C/O core leading to a thermonuclear supernova.
@@
The He donor channel as a mechanism to produce a double detonation was already investigated in the early 1980s by Taam (1980a, b), Nomoto (1982a, b). Later studies by Livne (1990), Livne & Glasner (1990), Langer (1991), Livne & Glasner (1991), Fink et al. (2007), and Fink et al. (2010) further explored He-accreting WDs and established that they are promising channels for SNe Ia. However, they also estimated that single degenerate double detonations require thick He shells which likely will lead to spectra representing a peculiar supernova Ia. The light curve of ZTF18aaqeasu (also known as SN 2018byg and ATLAS 18pqq), De et al. (2019), suggests that it is a double detonation supernova with a massive He shell ($\sim 0.15\,\mathrm{M_{\odot}}$). This thick He shell aligns with predictions by Bauer et al. (2017) for a hot subdwarf + WD binary (CD-30°11223). On the contrary, studies by Sim et al. (2010), Kromer et al. (2010), and Woosley & Kasen (2011) also explored channels that would enable the synthetic spectra to resemble a “normal” Type Ia supernova. Furthermore, Wang et al. (2013) and Neunteufel et al. (2016) probe the He star + WD binary channels as a potential single-degenerate double detonation mechanism. In addition, some studies explored the effects of WD rotation (Yoon & Langer, 2004) and magnetic fields (Neunteufel et al., 2017) on He ignition.
@@
He donor channels are also interesting as a mechanism to produce hypervelocity runway stars such as US 708. Should there be a full explosion of the WD, the hot subdwarf can be ejected at very high velocities (Justham et al., 2009; Wang & Han, 2009). This mechanism can explain some of the observed population of hypervelocity runaway stars with velocities of $\sim 1000\,\mathrm{km\,s^{-1}}$ (Geier et al., 2015; Brown et al., 2015; Neunteufel, 2020; Neunteufel et al., 2022).
@@
Given the increasing number of observed hot subdwarf and white dwarf binaries, interest in understanding the fate of these systems has been growing. Our galaxy contains at least $10^{3}-10^{4}$ hot subdwarf + WD binaries that are likely to interact within the subdwarf lifetime (Bauer & Kupfer, 2021). Despite detailed modeling studies of individual systems, the literature still lacks a broad study of binary spatial configurations that lead to double detonation supernovae. The purpose of this paper is to map the initial binary configurations of hot subdwarf and WD binaries to their fates: double detonation supernovae, He novae, and double white dwarfs.
@@ 2 Methods
Figure 2: An example of a binary with $\mathrm{M_{d}=0.5\,\mathrm{M_{\odot}}}$, $\mathrm{M_{a}=0.8\,\mathrm{M_{\odot}}}$, and $\mathrm{P_{i}=0.84\,hours}$, where a thermonuclear explosion occurs in the WD (accretor). The background colors yellow, blue, and green represent the gravitational-wave inspiral phase (GW-inspiral), H mass transfer phase (H-MT), and He mass transfer phase (He-MT), respectively. Panels (a) through (f) show the evolution of various parameters: (a) Mass Evolution: $M_{a}$ (accretor) and $M_{d}$ (donor); (b) Orbital Period Evolution; (c) Radius Evolution: $R_{d}$ (donor’s radius) and $R_{roche,d}$ (Roche radius of the donor); (d) mass transfer rate; (e) Evolution of surface mass fraction: $X_{H}$ (hydrogen) and $X_{He}$ (helium) in the donor; (f) He Mass on the Accretor; (g) Luminosity Evolution: $L_{a}$ (accretor), $L_{d}$ (donor), and $L_{\text{accretion}}$ (accretion). The accretor gains mass at the rate of $\sim 10^{-8}\,\mathrm{M_{\odot}yr^{-1}}$, resulting in the He ignition in the He layers denser than the assumed critical density for detonation ($>10^{6}\,\mathrm{g\,cm^{-3}}$).
@@
Figure 3: Each curve depicts the evolution of the temperature-density profile of the accreting white dwarf at various times, as indicated in the legend, for a system that undergoes double detonation supernova (left panel) and a He nova (right panel). The black star indicates the C/O-He boundary, that moves to higher density as the WD accretes and contracts. In the left panel, the WD accretes at the rate of $\sim 10^{-8}\,\mathrm{M_{\odot}yr^{-1}}$, igniting the high density regions of the He shell ($>10^{6}\,\mathrm{g\,cm^{-3}}$), and so we classify the outcome as a double detonation supernova. In the right panel, the accretor’s evolution is initially dominated by the cooling of the WD. After the accretion starts, the WD accretes at $\sim 10^{-7}\,\mathrm{M_{\odot}yr^{-1}}$. The WD experiences compressional heating, causing the low-density ($<10^{6}\,\mathrm{g\,cm^{-3}}$) layers to ignite explosively. We then classify this system as a He nova.
In this section, we describe the physical assumptions used in this study. We use Modules for Experiments in Stellar Astrophysics (modified00footnotetext: The composition of the accretion stream is modified to change the mass fraction of H to He throughout accretion. See Sec. 2.2 for more details. This is done by modifying the binary_mdot.f90 file. All files including the modified binary_mdot.f90 can be found at Zenodo. MESA version r23.05.1 Paxton et al. 2011, 2013, 2015, 2018, 2019; Jermyn et al. 2023) to model the binary evolution of hot subdwarf + WD binaries. The MESA Equation of state (EOS) is a blend of the OPAL (Rogers & Nayfonov, 2002), SCVH (Saumon et al., 1995), FreeEOS (Irwin, 2004), HELM (Timmes & Swesty, 2000), PC (Potekhin & Chabrier, 2010), and Skye (Jermyn et al., 2021) EOSes. Radiative opacities are primarily from OPAL (Iglesias & Rogers, 1993, 1996), with low-temperature data from Ferguson et al. (2005) and the high-temperature, Compton-scattering dominated regime by Poutanen (2017). Electron conduction opacities are from Cassisi et al. (2007) and Blouin et al. (2020). Nuclear reaction rates are from JINA REACLIB (Cyburt et al., 2010), NACRE (Angulo et al., 1999) and additional tabulated weak reaction rates Fuller et al. (1985); Oda et al. (1994); Langanke & Martínez-Pinedo (2000). Screening is included via the prescription of Chugunov et al. (2007). Thermal neutrino loss rates are from Itoh et al. (1996).
@@
We use Schwarzchild stability criterion to map the convective zones. Further, mixing by convection is followed using the mixing length theory, with mixing_length_alpha = 2 and MLT_option = ’TDC’, the time-dependent convection model from Kuhfuss (1986). To account for mass loss during the red giant branch, we use the Reimers prescription from MESA and switch to the T-Blöcker prescription during the asymptotic giant branch phase. We do not include the effects of convective overshooting while creating hot subdwarfs and WDs. We further assume no gravitational settling, meaning that the hot subdwarf/WD atmosphere contains not only hydrogen but also contamination from heavier elements. Nevertheless, the mass fraction of these elements is too low to affect the binary evolution or the final fate of the system.
In subsection 2.1, we explain the construction of individual models. Subsection 2.2 discusses the binary evolution setup. Finally, in subsection 2.3, we detail the conditions used to classify the systems as He nova and double detonation supernova.
2.1 Construction of models for hot subdwarf and WD
@@
Table 1: Initial parameters (masses and orbital periods) for hot subdwarfs and WDs. For more details, see Sec. 2.1.
A hot subdwarf star is a core He burning star with a very thin ($<10^{-2}\,\mathrm{M_{\odot}}$; Heber 2016) H envelope. To create starting models for hot subdwarfs, we first evolve single stars from the pre-main sequence until He ignition. The hot subdwarf gets stripped either during a mass transfer phase or a CE phase. In this work, we approximate the mass loss during the mass-transfer or CE phase by applying a high wind mass loss rate, while freezing the nuclear reactions in the core. The stripping continues until the total mass of H in the envelope is $3\times 10^{-4}\,\mathrm{M}_{\odot}$ [^1] This H mass is chosen to prevent thin-shell instabilities in the donor. In our models, hot subdwarf stars originating from low-mass progenitors ($<2.3\,\mathrm{M_{\odot}}$) have degenerate He cores, leading to He ignition via a core He flash. In contrast, hot subdwarf stars with progenitors $>$ $2.3\,\mathrm{M_{\odot}}$ ignite He gradually without undergoing a core He flash. The initial zero-age main sequence (ZAMS) masses of hot subdwarfs are chosen from the range $1\,\mathrm{M}_{\odot}$ to $5.95\,\mathrm{M}_{\odot}$ to include both degenerate and non-degenerate core He ignition. In our models, hot subdwarf stars are in the mass range of $0.33\,\mathrm{M}_{\odot}$ to $0.8\,\mathrm{M}_{\odot}$. The lower limit of $0.33\,\mathrm{M}_{\odot}$ represents the minimum mass required for core He ignition. The most massive model, $0.8\,\mathrm{M}_{\odot}$ expands to a giant-like phase reaching the radii of approximately $10\,\mathrm{R}_{\odot}$. Notably, our hot subdwarf models include $0.47\,\mathrm{M}_{\odot}$, which is considered the canonical mass for such stars (Han et al., 2002; Heber et al., 2003). Table. 1 gives the details of the donor masses.
We create carbon oxygen WDs (0.7 to $1.0\,\mathrm{M}_{\odot}$) after modifying the make_co_wd test suite to cool until the age of $3\times 10^{8}\,$ years. The modified test suite make_co_wd evolves a single star from the pre-main sequence phase to the WD phase. The mass range of the WD spans from 0.7 to $1.1\,\mathrm{M}_{\odot}$, chosen to align with the existing observations and to anticipate the potential explosive fate of these systems. The massive $1.1\,\mathrm{M}_{\odot}$ WD is created by rescaling the $1.0\,\mathrm{M}_{\odot}$ WD using the MESA inlist parameter relax_mass_scale, which rescales to the new mass without altering the composition profile. All the WD models are further refined to convert all H to 4He, avoiding the extra computational cost of evolving through classical novae at the onset of mass transfer, which is not the focus of this study.
2.2 Binary evolution
@@
Figure 4: An example of a binary with $\mathrm{M_{d}=0.75\,\mathrm{M_{\odot}}}$, $\mathrm{M_{a}=0.8\,\mathrm{M_{\odot}}}$, and $\mathrm{P_{i}=1.9\,hours}$, where a He nova occurs in the WD (accretor). The panels display the evolution of different parameters as in Fig. 2. The background colors yellow, and green represent the gravitational-wave inspiral phase (GW-inspiral), and mass transfer phase during the He shell burning phase (Shell-MT), respectively. The accretor accretes with a relatively high accretion rate of $\sim 10^{-7}\,\mathrm{M_{\odot}yr^{-1}}$, resulting in the ignition in the He layers less dense than the assumed critical density for detonation ($<10^{6}\,\mathrm{g\,cm^{-3}}$).
In this section, we explain the physical assumptions used in the binary evolution of hot subdwarf + white dwarf binaries and detail the chosen orbital period range.
When evolving the binary system, we use basic_plus_fe56_ni58.net for hot subdwarf stars and a custom nco.net from Bauer et al. 2017 (included in the Zenodo repository) for WDs. For the hot subdwarf, the nuclear network basic_plus_fe56_ni58.net contains stable isotopes ${}^{56}\text{Fe}$ and ${}^{58}\text{Ni}$ in addition to the basic isotopes ${}^{1}\text{H},\,^{3}\text{He},\,^{4}\text{He},\,^{12}\text{C},\,^{14}\text{N},\,^{16}\text{O},\,^{20}\text{Ne},\,^{24}\text{Mg}$. This has been chosen to account for all isotopes and reactions from He burning. For the WD, the nuclear network includes the NCO (14N$(e^{-},\nu)^{14}$$(\alpha,\gamma)^{18}$) chain, in which the electron capture to ${}^{14}\text{N}$ in high densities ($\sim 10^{6}\,\mathrm{g\,cm^{-3}}$), leads to the formation of ${}^{14}\text{C}$. This ${}^{14}\text{C}$ can further capture an alpha particle inducing a thermonuclear runaway in the accretor (Bauer et al., 2017).
For convection treatments in the He core of hot subdwarf, we employ “predictive mixing”, which allows the core to grow while avoiding breathing pulses during the late stages of He core burning. For details on convection in hot subdwarfs in general, see Ostrowski et al. (2021).
The mass transfer rate follows MESA’s Kolb scheme (Kolb & Ritter, 1990; Paxton et al., 2015). We assume conservative mass transfer for our models. Thus, the mass transfer rate is always equal to the accretion rate. The entropy of the accreted material is determined by matching the entropy of the outer shell of the accretor. Further, the WDs undergoing accretion are expected to undergo compressional heating, which MESA incorporates into evolutionary models through its energy equation and thermodynamically consistent EOS (see Jermyn et al. 2023). Exploring the effects on non-conservative mass transfer, tidal heating, and spin-up of the accretor is left to future work.
The donor starts transferring the envelope, which adds H and heats the surface of the WD, resulting in classical novae. Simulating and evolving through these novae is computationally expensive and not the focus of this paper. Therefore, we modify the accretion stream to convert the mass fraction of H and ${}^{3}\text{He}$ to the mass fraction of ${}^{4}\text{He}$. Furthermore, the initial composition profile of the WD model is further modified to convert ${}^{3}\text{He}$ to ${}^{4}\text{He}$ to avoid further production of H through nuclear reactions.
We initiate our hot subdwarf and WD models in a wide range of periods starting from 36 minutes to 7.2 hours. This range is chosen to encompass the mass transfer during various core He fractions for all our donors, including phases like shell burning phase, and thermal pulses. Additionally, the period range allows for donors to evolve without interaction to become a WD. Further details of the grid can be found in the Table 1.
The evolution is followed until either the accretor reaches the ignition conditions (see Sec. 2.3) or the donor evolves to a WD making a double white dwarf. We terminate our simulations as double white dwarfs if the binary system is non-interacting and the radius of the hot subdwarf decreases to $0.03\,\mathrm{R_{\odot}}$. Once this radius is attained, all the hot subdwarf-turned WDs in our simulations enter the cooling track.

### v2

@@
Abinaya Swaruba Rajamuthukumar    Evan B. Bauer    Stephen Justham    Rüdiger Pakmor
Selma E. de Mink
   Patrick Neunteufel
Compact binaries containing hot subdwarfs and white dwarfs have the potential to evolve into a variety of explosive transients. These systems could also explain hypervelocity runaway stars such as US 708. We use the detailed binary evolution code MESA to evolve hot subdwarf and white dwarf stars interacting in binaries. We explore their evolution toward double detonation supernovae, helium novae, or double white dwarfs. We present a grid of 3120 binary evolution models that map from initial conditions, such as the orbital period and masses of the hot subdwarf and white dwarf, to these outcomes. The minimum amount of helium required to ignite the helium shell that leads to a double detonation supernova in our grid is $\approx 0.05\,\mathrm{M_{\odot}}$, likely too large to produce spectra similar to normal Type Ia supernovae, but compatible with inferred helium shell masses from some observed peculiar Type I supernovae. We also provide the helium shell masses for our double white dwarf systems, with a maximum He shell mass of $\approx 0.18\,\mathrm{M_{\odot}}$. In our double detonation systems, the orbital velocity of the surviving donor star ranges from $\approx 450\,\mathrm{km\,s^{-1}}$ to $\approx 1000\,\mathrm{km\,s^{-1}}$. Among the surviving donors, we also estimate the runaway velocities of proto-white dwarfs, which have higher runaway velocities than hot subdwarf stars of the same mass. Our grid will provide a first-order estimate of the potential outcomes for the observation of binaries containing hot subdwarfs and white dwarfs from future missions like Gaia, LSST, and LISA.
@@ 1 Introduction
Hot subdwarfs are stripped helium (He) stars that burn He in their core with little to no hydrogen (H) envelope (1986A&A...155...33H; 2016PASP..128h2001H). Studies such as 2002MNRAS.336..449H; 2003MNRAS.341..669H suggest that binary interactions such as stable or unstable mass transfer plays a major role in the formation of hot subdwarfs. Observations provide evidence that most of the hot subdwarfs are found with binary companions (e.g., 2020A&A...642A.180P). Further, a significant fraction of hot subdwarfs are found in close orbits ($P_{\mathrm{orb}}\sim 10\,\text{days}$) with white dwarfs (WDs), which are likely post common envelope systems (2002MNRAS.336..449H; 2003MNRAS.341..669H).
@@
Figure 1: Schematic diagram representing possible evolution channels of a hot subdwarf + WD binary to a double white dwarf, a double detonation supernova or a He nova. Double detonation supernovae and He novae result from systems that undergo mass transfer and later experience thermonuclear instability in the He layer on the surface of the WD. The differentiation between double detonation supernova and He nova is based on the He layer’s critical density (see Sect. 2.1). Double white dwarfs result from systems that either evolve in isolation or involve episodes of accretion but do not enter unstable nuclear burning.
@@
Compact binaries containing hot subdwarfs and WDs have gained significant interest following the discovery of a super-Chandrasekhar binary by 2000MNRAS.317L..41M. This binary is particularly interesting as a potential Type Ia supernova progenitor (but, see also 2001A&A...376L...9E for other outcomes). The number of discoveries of binaries containing hot subdwarfs with WD companions has steadily increased (e.g., 2004Ap&SS.291..321N; 2012ApJ...759L..25V; Geier2013; 2023MNRAS.525..183S). Large spectroscopic studies have been characterizing binaries containing hot subdwarfs (e.g., 2022A&A... [^6] A.182S). In addition to those detached binaries, 2020ApJ...898L..25K; 2020ApJ...891...45K identified binaries in which the hot subdwarf is Roche-lobe filling. Furthermore, some of recent the observed binaries such as CD-30°11223 (2012ApJ...759L..25V; Geier2013) and HD 265435 (2021NatAs...5.1052P) are verification binaries for LISA (for more details see 2024ApJ...963..100K) and suggested progenitors of double donation supernovae (2022ApJ...925L..12K).
@@
Interacting hot subdwarf + WD binaries could potentially result either in an explosion only in the He shell of the WD (a He nova, 2000IAUC.7552....1K; 2010ApJ...720..581G; 2014MNRAS.445.3239P) or a thermonuclear explosion of a WD (probably as a supernova Type Ia, 1973ApJ...186.1007W; 1984ApJS...54..335I). The thermonuclear explosion of the WD could result from a Chandrasekhar mass explosion (2003A&A...412L..53Y; 2009MNRAS.395..847W; 2016ApJ...821...28B) or from a sub-Chandrasekhar mass double detonation explosion due to He accretion. In the double detonation mechanism, He-accreting WDs undergo the first detonation in the He shell that triggers the second detonation in the C/O core leading to a thermonuclear supernova.
@@
The He donor channel as a mechanism to produce a double detonation was already investigated in the early 1980s by 1980ApJ...237..142T; 1980ApJ...242..749T, 1982ApJ...253..798N; 1982ApJ...257..780N. Later studies by 1990ApJ...354L..53L, 1990ApJ...361..244L, 1991A&A...252..669L, 1991ApJ...370..272L, 2007A&A...476.1133F, and 2010A&A...514A..53F further explored He-accreting WDs and established that they are promising channels for SNe Ia. However, they also estimated that single degenerate double detonations require thick He shells which likely will lead to spectra representing a peculiar supernova Ia. The light curve of ZTF18aaqeasu (also known as SN 2018byg and ATLAS 18pqq), 2019ApJ...873L..18D, suggests that it is a double detonation supernova with a massive He shell ($\sim 0.15\,\mathrm{M_{\odot}}$). This thick He shell aligns with predictions by 2017ApJ...845...97B for a hot subdwarf + WD binary (CD-30°11223). On the contrary, studies by 2010ApJ...714L..52S, 2010ApJ...719.1067K, and 2011ApJ...734...38W also explored channels that would enable the synthetic spectra to resemble a “normal” Type Ia supernova. Furthermore, 2013A&A...559A..94W and 2016A&A...589A..43N probe the He star + WD binary channels as a potential single-degenerate double detonation mechanism. In addition, some studies explored the effects of WD rotation (2004A&A...419..623Y) and magnetic fields (2017A&A...602A..55N) on He ignition.
@@
He donor channels are also interesting as a mechanism to produce hypervelocity runway stars such as US 708. Should there be a full explosion of the WD, the donor can be ejected at very high velocities (2009A&A...493.1081J; 2009A&A...508L..27W). Depending on the time of explosion of the accretor the donor star can either be a He star or a white dwarf. This mechanism could explain some of the observed population of hypervelocity runaway stars with velocities of $\lesssim 1000\,\mathrm{km\,s^{-1}}$(2015Sci...347.1126G; 2015ApJ...804...49B; 2020A&A...641A..52N; 2022A&A...663A..91N).
@@
Given the increasing number of observed hot subdwarf and white dwarf binaries, interest in understanding the fate of these systems has been growing. Our galaxy contains at least $10^{3}-10^{4}$ hot subdwarf + WD binaries that are likely to interact within the subdwarf lifetime (2021ApJ...922..245B). Despite detailed modeling studies of individual systems, the literature still lacks a broad study of binary spatial configurations that lead to double detonation supernovae. The purpose of this paper is to map the initial binary configurations of hot subdwarf and WD binaries to their fates: double detonation supernovae, He novae, and double white dwarfs.
@@ 2 Methods
Figure 2: Example of a binary with $\mathrm{M_{d}=0.5\,\mathrm{M_{\odot}}}$, $\mathrm{M_{a}=0.8\,\mathrm{M_{\odot}}}$, and $\mathrm{P_{i}=0.84\,hours}$, where a thermonuclear explosion occurs in the WD (accretor). The background colors yellow, blue, and green represent the gravitational-wave inspiral phase (GW-inspiral), H mass transfer phase (H-MT), and He mass transfer phase (He-MT), respectively. Panels (a) through (f) show the evolution of various parameters: (a) Mass Evolution: $M_{a}$ (accretor) and $M_{d}$ (donor); (b) Orbital Period Evolution; (c) Radius Evolution: $R_{d}$ (donor’s radius) and $R_{roche,d}$ (Roche radius of the donor); (d) mass transfer rate; (e) Evolution of surface mass fraction: $X_{H}$ (hydrogen) and $X_{He}$ (helium) in the donor; (f) He Mass on the Accretor; (g) Luminosity Evolution: $L_{a}$ (accretor), $L_{d}$ (donor), and $L_{\text{accretion}}$ (accretion). The accretor gains mass at the rate of $\sim 10^{-8}\,\mathrm{M_{\odot}yr^{-1}}$, resulting in the He ignition in the He layers denser than the assumed critical density for detonation ($10^{6}\,\mathrm{g\,cm^{-3}}$).
@@
Figure 3: Each curve depicts the evolution of the temperature-density profile of the accreting white dwarf at various times, as indicated in the legend, for a system that undergoes double detonation supernova (left panel) and a He nova (right panel). The black star indicates the C/O-He boundary, that moves to higher density as the WD accretes and contracts. In the left panel, the WD accretes at the rate of $\sim 10^{-8}\,\mathrm{M_{\odot}yr^{-1}}$, igniting the high density regions of the He shell ($10^{6}\,\mathrm{g\,cm^{-3}}$), and so we classify the outcome as a double detonation supernova. In the right panel, the accretor’s evolution is initially dominated by the cooling of the WD. After the accretion starts, the WD accretes at $\sim 10^{-7}\,\mathrm{M_{\odot}yr^{-1}}$. The WD experiences compressional heating, causing the low-density ($10^{6}\,\mathrm{g\,cm^{-3}}$) layers to ignite explosively. We then classify this system as a He nova.
In this section, we describe the physical assumptions used in this study. We use Modules for Experiments in Stellar Astrophysics (modified [^1] The composition of the accretion stream is modified to change the mass fraction of H to He throughout accretion. See later in Sect. 2 for more details. This is done by modifying the binary_mdot.f90 file. All files including the modified binary_mdot.f90 can be found at Zenodo. MESA version r23.05.1 Paxton2011; Paxton2013; Paxton2015; 2018ApJS..234...34P; Paxton2019; 2023ApJS..265...15J) to model the binary evolution of hot subdwarf + WD binaries.
Convective Treatment: We use the Schwarzschild stability criterion to determine the regions that are unstable to convection. Further, mixing by convection is followed using the mixing length theory (1958ZA.....46..108B), with mixing_length_alpha = 2 and MLT_option = ’TDC’, the time-dependent convection model based on 1986A&A...160..116K. For convection treatments in the He core of hot subdwarf, we employ “predictive mixing,” which allows the core to grow while avoiding breathing pulses during the late stages of He core burning. For details on convection in hot subdwarfs in general, see 2021MNRAS.503.4646O.
[…]
Stellar winds: To account for mass loss during the red giant branch, we use the Reimers prescription (1975MSRSL...8..369R) from MESA and switch to the Blocker prescription (1995A&A...299..755B) during the asymptotic giant branch phase. Further, we assume no winds during the hot subdwarf phase. The mass and luminosity of these low-mass stripped stars are too low to drive strong winds that would affect the binary evolution and hence the results (2016A&A...593A.101K).
Hot subdwarf: A hot subdwarf star is a low mass core He burning star with a very thin ($\lesssim 10^{-2}\,\mathrm{M_{\odot}}$; 2016PASP..128h2001H) H envelope. To create starting models for hot subdwarfs, we first evolve single stars from the pre-main sequence until He ignition. The hot subdwarf gets stripped either during a mass transfer phase or a CE phase. In this work, we approximate the mass loss during the mass-transfer or CE phase by applying a high wind mass loss rate, while freezing the nuclear reactions in the core. The stripping continues until the total mass of H in the envelope is $3\times 10^{-4}\,\mathrm{M}_{\odot}$ [^2] This H mass is chosen to prevent thin-shell instabilities in the donor. In our models, hot subdwarf stars originating from low-mass progenitors ($2.3\,\mathrm{M_{\odot}}$) have degenerate He cores, leading to He ignition via a core He flash. In contrast, hot subdwarf stars with progenitors >> $2.3\,\mathrm{M_{\odot}}$ ignite He gradually without undergoing a core He flash. The initial zero-age main sequence (ZAMS) masses of hot subdwarfs are chosen from the range $1\,\mathrm{M}_{\odot}$ to $5.95\,\mathrm{M}_{\odot}$ to include both degenerate and non-degenerate core He ignition. The hot subdwarf stars in our models are in the mass range of $0.33\,\mathrm{M}_{\odot}$ to $0.8\,\mathrm{M}_{\odot}$. The lower limit of $0.33\,\mathrm{M}_{\odot}$ represents the minimum mass required for core He ignition. The most massive model, $0.8\,\mathrm{M}_{\odot}$ expands to a giant-like phase reaching the radii of approximately $10\,\mathrm{R}_{\odot}$. Notably, our hot subdwarf models include $0.47\,\mathrm{M}_{\odot}$, which is considered the canonical mass for such stars (2002MNRAS.336..449H; 2003A&A...411L.477H). Table. 1 gives the details of the donor masses.
White dwarf: We create carbon oxygen WDs (0.7 to $1.0\,\mathrm{M}_{\odot}$) after modifying the make_co_wd test suite to cool until the age of $3\times 10^{8}\,$ years. The modified test suite make_co_wd evolves a single star from the pre-main sequence phase to the WD phase. The mass range of the WD spans from 0.7 to $1.1\,\mathrm{M}_{\odot}$, chosen to align with the existing observations and to anticipate the potential explosive fate of these systems. The massive $1.1\,\mathrm{M}_{\odot}$ WD is created by rescaling the $1.0\,\mathrm{M}_{\odot}$ WD using the MESA inlist parameter relax_mass_scale, which rescales to the new mass without altering the composition profile. All the WD models are further refined to convert all H to 4He, avoiding the extra computational cost of evolving through classical novae at the onset of mass transfer, which is not the focus of this study.
Binary Evolution: The mass transfer rate follows MESA’s Kolb scheme (1990A&A...236..385K; Paxton2015). We assume conservative mass transfer for our models. Thus, the mass transfer rate is always equal to the accretion rate. The entropy of the accreted material is determined by matching the entropy of the outer shell of the accretor. Further, the WDs undergoing accretion are expected to undergo compressional heating, which MESA incorporates into evolutionary models through its energy equation and thermodynamically consistent EOS (see 2023ApJS..265...15J). Exploring the effects on non-conservative mass transfer, tidal heating, and spin-up of the accretor is left to future work.
The donor starts transferring the envelope, which adds H and heats the surface of the WD, resulting in classical novae. Simulating and evolving through these novae is computationally expensive and not the focus of this paper. Therefore, we modify the accretion stream to convert the mass fraction of H and $^{3}\text{He}$ to the mass fraction of $^{4}\text{He}$. This treatment is applied consistently throughout the mass transfer phase. Furthermore, the initial composition profile of the WD model is further modified to convert $^{3}\text{He}$ to $^{4}\text{He}$ to avoid further production of H through nuclear reactions.
We initiate our hot subdwarf and white dwarf models on a regular orbital period grid spanning 36 minutes to 7.2 hours, excluding systems that are Roche-lobe filling at the onset. This range captures mass transfer at various core He fractions for the donor stars, including phases such as shell burning and thermal pulses. Additionally, the period range allows for donors to evolve without interaction to become a WD. Further details of the grid can be found in the Table 1.
The evolution is followed until either the accretor reaches the ignition conditions (see Sect. 2.1) or the donor evolves to a WD making a double white dwarf. We terminate our simulations as double white dwarfs if the binary system is non-interacting and the radius of the hot subdwarf decreases to $0.03\,\mathrm{R_{\odot}}$. Once this radius is attained, all the hot subdwarf-turned WDs in our simulations enter the cooling track.
@@
We do not include the effects of convective overshooting while creating hot subdwarfs and WDs. All our models assume $Z=0.02$, with the initial metal fractions taken from Grevesse1998. We further assume no gravitational settling, meaning that the hot subdwarf or white dwarf atmosphere contains not only hydrogen but also contamination from heavier elements. Nevertheless, the mass fraction of these elements is likely too low to significantly impact the binary evolution or the final fate of the system.
@@
Table 1: Initial parameters (masses and orbital periods) for hot subdwarfs and WDs. For more details, see Sect. 2.
@@
Figure 4: Example of a binary with $\mathrm{M_{d}=0.75\,\mathrm{M_{\odot}}}$, $\mathrm{M_{a}=0.8\,\mathrm{M_{\odot}}}$, and $\mathrm{P_{i}=1.9\,hours}$, where a He nova occurs in the WD (accretor). The panels display the evolution of different parameters as in Fig. 2. The background colors yellow, and green represent the gravitational-wave inspiral phase (GW-inspiral), and mass transfer phase during the He shell burning phase (Shell-MT), respectively. The accretor accretes with a relatively high accretion rate of $\sim 10^{-7}\,\mathrm{M_{\odot}yr^{-1}}$, resulting in the ignition in the He layers less dense than the assumed critical density for detonation ($10^{6}\,\mathrm{g\,cm^{-3}}$).